40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 6 02:09:39 2011, rana, Update, Electronics, POY11 Rework Nearly Complete 777.png
    Reply  Mon Jun 6 16:44:34 2011, Koji, Update, Electronics, POY11 tested POY11_test.pdf
Message ID: 4786     Entry time: Mon Jun 6 02:09:39 2011     Reply to this: 4787
Author: rana 
Type: Update 
Category: Electronics 
Subject: POY11 Rework Nearly Complete 

I've finished tuning POY11 and it is now sitting on top of the analyzer waiting for Koji to test its noise.

 Notes:

  1. R2 has been switched from a 50 -> 110 Ohm R and a (29 Ohm | 47 pF) in parallel to it. This makes the gain of the MAX4107 be ~20 above 100 MHz and ~5 around 11 MHz. The high gain at high frequency makes it stable (squashes the 200 MHz oscillation) and the low gain at low frequency is to prevent saturation (the raw 11 MHz transimpedance is too high).
  2. There are two notches: 22.12 MHz and 55.3 MHz.
  3. The TEST IN input has been disabled since it wasn't useful.
  4. 107 Ohms inserted into between the U11 output and the TSENSE feedthrough. This is to prevent oscillations when driving the long Dsub cable. This needs to be done on all Gold Box RFPDs.
  5. R4 -> 50 Ohms.
  6. Had trouble tuning this to get the resonance at 11.06 MHz. This turned out to be the parallel inductance coming from L3 (previously 1.45 uH) whereas we needed a total inductance of ~1.6 uH. So I changed L3 to 33 uH to get it to be negligible compared to 1.7 uH at 11 MHz. This needs to be considered for all the 11 MHz diodes.

 

Attachment 1: 777.png  10 kB  | Hide | Hide all
777.png
ELOG V3.1.3-